OpenStudy (anonymous):

PLEASE HELP PLEASE!does anyone know how find a solution to a system of equations such as: 2y=-3x+y=17 3x+y=25

OpenStudy (anonymous):

2y=-3x+y=17 is this right?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

we will Re-arrange it

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

-3x+y-2y=17 3x+y=25

OpenStudy (anonymous):

Now we will add the two equations

OpenStudy (anonymous):

ok and how do u do tht?

OpenStudy (anonymous):

-3x-y=17 3x+y=25

OpenStudy (anonymous):

i am confused now

OpenStudy (anonymous):

+3x and -3x cancel +y and -y cancel

OpenStudy (anonymous):

ahhhh ok

OpenStudy (anonymous):

0=42(WHICH IS WRONG) so your system is no valid

OpenStudy (anonymous):

isnt it 8,1?

OpenStudy (anonymous):

how, show me what you did

OpenStudy (anonymous):

ok so i got -3x-y=25 rearranging the problem and then i subtracted and plugged

OpenStudy (anonymous):

how did you get -3x-y=25 when original equation is 3x+y=25

OpenStudy (anonymous):

can i give you a different problem because i wanted to check if 2y=-3+6 and-9-6y=18 ok so rearranged to 2y=-3y+6 and 6y=18-6y does that sound good so far?

OpenStudy (anonymous):

Your orginial Rearrange 2y=-3+6 2y=-3y+6 -9-6y=18 6y=18-6y

OpenStudy (anonymous):

ok then wht do u do after tht and i think for the rearranged bottom one should it be 9x instead of 9y cause i typed the wrong data sorry

OpenStudy (anonymous):

and the 3 for the top should have an x instead of a y

OpenStudy (anonymous):

im just confused on what to do AFTER rearranging the equations

OpenStudy (anonymous):

If 2y=-3x+y=17 is the correct problem, then you can extract 3 equations from that : 2y=-3x+y -3x+y=17 2y=17 The 1st equation simplifies to y=-3x. Substituting that into the 2nd equation gives us -3x+-3x=17 which simplies to -6x=17. Solving that for x gives us x=-17/6. From the 3rd equation 2y=17 we get that y=17/2. So we have as the solution x=-17/6 and y=17/2